### v3

@@ Abstract
People acquire concepts through rich physical and social experiences and use them to understand the world. In contrast, large language models (LLMs), trained exclusively through next-token prediction over language data, exhibit remarkably human-like behaviors. Are these models developing concepts akin to humans, and if so, how are such concepts represented and organized? To address these questions, we reframed the classic reverse dictionary task to simulate human concept inference in context and investigated the emergence of human-like conceptual representations within LLMs. Our results demonstrate that LLMs can flexibly derive concepts from linguistic descriptions in relation to contextual cues about other concepts. The derived representations converged towards a shared, context-independent structure that effectively predicted human behavior across key psychological phenomena, including computation of similarities, categories and semantic scales. Moreover, these representations aligned well with neural activity patterns in the human brain, even in response to visual rather than linguistic stimuli, providing evidence for biological plausibility. These findings establish that structured, human-like conceptual representations can naturally emerge from language prediction without real-world grounding. More broadly, our work positions LLMs as promising computational tools for understanding complex human cognition and paves the way for better alignment between artificial and human intelligence.
@@ 1 Introduction
Humans are able to construct mental models of the world and use them to understand and navigate their environment [1, 2]. Central to this ability is the capacity to form broad concepts that constitute the building blocks of these models [3]. These concepts, often regarded as mental representations, capture regularities in the world while abstracting away extraneous details, enabling flexible generalization to novel situations [4]. For example, the concept of sun can be instantly formed and deployed across diverse contexts: observing it rise or set in the sky, yearning for its warmth on a chilly winter day, or encountering someone who exudes positivity. Its role in the solar system can be analogized to the nucleus in an atom, enriching learning and understanding. The nature of concepts has long been a focus of inquiry across philosophy, cognitive science, neuroscience, and linguistics [5, 6, 7, 8, 9, 10, 11, 12]. These investigations have uncovered diverse properties that concepts need to satisfy, often framed by the long-standing divide between symbolism and connectionism. Symbolism emphasizes discrete, explicit symbols with structured and compositional properties, enabling abstract reasoning and recombination of ideas [13, 14]. In contrast, connectionism conceptualizes concepts as distributed, emergent patterns across networks, prioritizing continuity and gradedness, which excel in handling noisy inputs and learning from experience [15, 16]. Although there is growing consensus on the need to integrate the strengths of both paradigms to account for the complexity and richness of human concepts [17, 18], reconciling the competing demands remains a significant challenge.
@@
Recent advances in large language models (LLMs) within artificial intelligence (AI) have exhibited human-like behavior across various cognitive and linguistic tasks, from language generation [19, 20] to decision-making [21] and reasoning [22, 23, 24]. It has sparked intense interest and debate about whether these models are approaching human-like cognitive capacities based on concepts [25, 26, 27, 28, 29, 30]. Some argue that LLMs, trained exclusively on text data for next-token prediction, operate only on statistical associations between linguistic forms and lack concept-based understanding grounded in physical and social situations [31, 32, 29]. This argument is evidenced by their inconsistent performance, which often reveals vulnerabilities such as non-human-like errors and over-sensitivity to minor contextual variations [21, 33, 34]. Conversely, others contend that the performance of a system alone is insufficient to characterize its underlying competence [35], and the extent to which concepts should be grounded remains open [36, 37]. Instead, language may provide essential cues for people’s use of concepts and enable LLMs to approximate fundamental aspects of human cognition, where meaning arises from the relationships among concepts [38, 39]. Despite the conflicting views, there is broad consensus on the central role of concepts in human cognition. The core questions driving the debates are whether LLMs possess human-like conceptual representations and organization and, consequently, whether they can illuminate the nature of human concepts.
@@
To address these questions, we investigated the emergence of human-like conceptual representations from language prediction in LLMs. Our approach unfolded in three stages. First, we examined LLMs’ capacities to derive conceptual representations, focusing on the definitional and structured properties of human concepts. We leveraged LLMs’ in-context learning ability and guided them to infer concepts from descriptions with a few contextual demonstrations (Fig. 1). The models’ outputs offered a behavioral lens to evaluate the derived representations. We explored the organization of these representations by analyzing how their relational structures varied across different contextual demonstrations. Second, we assessed how well the LLM-derived representational structures align with psychological measures and investigated whether they capture rich, context-dependent human knowledge. These questions were tested by leveraging computations over the representations to predict human behavioral judgments. Finally, we mapped the LLM-derived conceptual representations to neural activity patterns in the human brain and analyzed their biological plausibility. Our experiments spanned thousands of naturalistic object concepts and beyond [40], providing a comprehensive analysis of LLMs’ conceptual representations. The findings reveal that language prediction can give rise to human-like conceptual representations and organization in LLMs. These representations integrated many valued aspects of previous accounts, combining the definitional and structural focus of symbolism with the continuity and graded nature of connectionist models. This work underscores the profound connection between language and human conceptual structures and demonstrates that LLM-derived conceptual representations offer a promising foundation for understanding human concepts.
@@ 2.1 Reverse dictionary as a conceptual probe
We reframed the reverse dictionary task as a domain-general approach to probe LLMs’ capacity for concept inference. A reverse dictionary identifies words based on provided definitions or descriptions [41], which is simple yet relevant to concept understanding and usage. For instance, consider a child learning that the concept moon corresponds to both “a round, glowing object that hangs up in the sky at night” and “Earth’s natural satellite”. The child then uses the term “moon”, rather than “circle” or “star”, to refer to this concept. The shared term “moon” helps interlocutors gauge alignment in their understanding of the concept. Instead of relying on a single, potentially ambiguous word (e.g., “bat”), the word-retrieval paradigm combines the words in descriptions to construct coherent meaning, inferring the corresponding concepts and mapping them back to words. The “form-meaning-form” process offers a targeted probe into the models’ capacity to contextually form meaningful representations and appropriately refer to them in ways that align with human understanding.
To guide LLMs through the process, we took advantage of their in-context learning ability and presented them with a small number of demonstrations in a reverse-dictionary format (“$[\textrm{Description}]\Rightarrow[\textrm{Word}]$”), followed by a query description (Fig. 1). We then compared model-generated completions to the intended term of the concept corresponding to the query description, thereby evaluating how well the models derived conceptual representations that aligned with human understanding. Formally, given $N$ pairs of descriptions and words as contextual demonstrations, an LLM $\mathcal{M}$ encodes the query description of a concept $c$ into a representation $\mathbf{h}_{c}=\textrm{encode}_{\mathcal{M}}\left(\textrm{description}\mid\textrm{context}_{N}\right)$, and generates a term based on it: $\textrm{predict}_{\mathcal{M}}\left(\mathbf{h}_{c}\right)$. The term was then compared to the name of the query concept shared by humans: $\textrm{name}_{\mathcal{H}}\left(c\right)$. We took $\mathbf{h}_{c}$ as the representation for concept $c$, which immediately precedes the following term that should be in semantic correspondence to the query description. Importantly, the demonstrations provided minimal and controllable context, and the query description served as a special case of general input for concept inference. We can analyze how models construct conceptual representations in response to different contextual cues by varying the number of demonstrations and the specific description-word pairings within them.
Figure 1: Illustration of the reverse dictionary task as a conceptual probe. A Transformer-based LLM is presented with $N$ description-word pairs as demonstrations in context, followed by a query description. The model is then prompted to encode the query description into a conceptual representation and predict the term that best matches the described concept.
@@ 2.2 Deriving concepts from definitional descriptions
We investigated whether LLMs can construct concepts from definitional descriptions through the reverse dictionary task. Leveraging data from the THINGS database [42], we prompted LLMs with randomly selected description-word pair demonstrations and examined their capacity to predict appropriate terms for the query descriptions (Methods 4.2). Fig. 2 shows the performance of LLaMA3-70B, a state-of-the-art open-source LLM, averaged across five independent runs at varying numbers of demonstrations. The exact match accuracy improved progressively as the number of demonstrations increased from 1 to 24, rising from $79.51\%$ (±$\pm 0.30\%$) to $89.45\%$ (±$\pm 0.30\%$). Gains were marginal beyond this threshold. These results indicate that, with a few demonstrations as context, LLMs can effectively combine words into coherent representations and reliably infer the corresponding concepts, despite varying contextual cues.
@@
Figure 2: Performance of LLaMA3-70B on the reverse dictionary task measured through exact match accuracy. Across the four plots, the black lines denote the model’s performance when presented with $N$ demonstrations sampled from the training set and evaluated on an independent test set. The blue and red lines illustrate the model’s behavior when presented with one misleading demonstration—a description paired with a proxy label (as specified in the figure titles)—and $N-1$ correct demonstrations of other concepts. The blue line shows the frequency with which the model copies the misleading demonstration and reproduces the proxy label, while the red line indicates how often it generates the correct word for the query concept given contextual information from other concepts. Shaded areas denote 95% bootstrapped confidence intervals, calculated from 10,000 resamples over five independent runs.
Furthermore, we extended our analyses to a broader range of descriptions and concepts to assess the generality of our findings derived from the THINGS database. Using data from WordNet, our extended analysis demonstrated the strong adaptability of LLMs across concepts differing in concreteness and word classes (Fig. S8a). Consistent results were also observed across various descriptions of the same concept (Fig. S8b). Introducing varying degrees of word order permutations to the query descriptions revealed that LLM predictions were sensitive to linguistic structure degradation when combining words to form concepts (Fig. S8c). These findings highlight the model’s effectiveness in concept inference and its ability to capture at least some of the compositional structure in natural language. Beyond LLaMA3-70B, we tested 66 additional open-source LLMs with different model architectures, scales and training data. The results revealed trends similar to those observed with LLaMA3-70B. They also indicated that larger models generally perform better and more effectively leverage contextual cues for concept inference (Supplementary Information SI 1.1, Fig. S8d–f).
@@ 2.3 Uncovering a context-independent conceptual structure
To uncover how concepts are represented and organized within LLMs, we looked into the representation spaces they constructed for concept inference and analyzed the interrelationships among the conceptual representations. We characterized the representation spaces formed under different contexts by their relational structure, captured through a (dis)similarity matrix, and measured pairwise alignment by correlating these matrices [43, 44] (Methods 4.3). For concepts in the THINGS database, the alignment between representation spaces gradually increased as the number of demonstrations rose from 1 to 24, with diminishing gains beyond this threshold (Fig. 3a). The alignment with the space formed by 120 demonstrations increased from $0.800$ (±$\pm 0.017$) with one demonstration to $0.970$ (±$\pm 0.003$) after 24 demonstrations. This alignment demonstrated a strong correlation with the model’s exact match accuracy on concept inference ($\rho=0.976$, $P<0.0001$, $95\%$ CI: 0.7300.7300.7300.7301.01.01.01.0) (Fig. 3b). These results suggest that LLMs are able to construct a coherent and context-independent relational structure, which is reflected in their concept inference capacity. Additional metrics corroborated this conclusion (Methods 4.3 and Supplementary Information SI 1.2), highlighting that intricate relationships among concepts are consistently preserved within LLMs’ representation spaces across varying contexts. This invariance supports the generalization of knowledge encoded in these relationships.
To examine whether different LLMs trained for language prediction can develop similar conceptual representations, we compared the representation spaces they formed based on 24 demonstrations. Using t-SNE [45], we visualized pairwise alignments between models and observed that LLMs with over $70\%$ exact match accuracy on concept inference clustered closely, whereas those with accuracy below $50\%$ exhibited greater dispersion (Fig. 3c). This indicates that better-performing models share more similar relational structure of concepts, while those with weak performance diverge in their own ways. Further supporting this observation, a correlational analysis demonstrated that models with higher exact match accuracy aligned more closely with representations derived from LLaMA3-70B ($\rho=0.870$, $P<0.0001$, $95\%$ CI: 0.7560.7560.7560.7560.9400.9400.9400.940; Fig. 3d). Additionally, when quantifying model complexity based on scale and training data (Methods 4.4), we found that higher-complexity LLMs tended to align better with LLaMA3-70B, though exceptions likely stemmed from constraints in model scale or training data quality (Fig. 3d). These findings suggest that, with sufficient model scale and extensive, high-quality training data, LLMs can converge towards a shared conceptual structure. This convergence, reflected in their concept inference capacity, arises naturally from language prediction without real-world reference.
@@
Figure 3: LLMs converge toward a similar representational structure of concepts. a, Alignment correlation (RSA) between the LLM-derived conceptual representations across different contextual demonstrations. The LLM (LLaMA3-70B) is presented with $N$ demonstrations, repeated across five independent runs. b, LLM performance on the reverse dictionary task reflects alignment with the representations formed based on 120 demonstrations. Each point corresponds to the representations formed by the LLM based on $N$ demonstrations, with the x-axis indicating exact match accuracy and the y-axis denoting the alignment correlation. Error bars indicate $95\%$ confidence intervals, calculated using 10,000 bootstrap resamples from five independent runs. c, A t-SNE visualization of conceptual representations produced by different LLMs. Each point corresponds to an LLM, with representations formed based on 24 demonstrations. Distances between them are calculated as $1-\textrm{alignment}$, with results averaged over five independent runs. LLMs are plotted in proportion to their complexity and color-coded by their exact match accuracy on the reverse dictionary task. Better-performing models (blue) exhibit more aligned representations. d, LLMs’ concept inference capacity reflects alignment correlation with conceptual representations derived from LLaMA3-70B. The x-axis represents the average performance on the reverse dictionary task with 24 demonstrations, while the y-axis indicates the alignment correlation with LLaMA3-70B representations. Data points correspond to conceptual representations derived from different models and are color-coded by model complexity. Linear fits are shown as straight lines, with shaded areas representing $95\%$ confidence intervals derived from 10,000 bootstrap resamples. Overall, models with higher complexity align more closely with the reference. An exception is Qwen2-0.5B, which, despite exhibiting relatively high complexity, has a low alignment score. This may be due to constraints in model scale and the quality of training data.

### v4

@@ Abstract
People acquire concepts through rich physical and social experiences and use them to understand and navigate the world. In contrast, large language models (LLMs), trained solely through next-token prediction on text, exhibit strikingly human-like behaviors. Are these models developing concepts akin to those of humans? If so, how are such concepts represented, organized, and related to behavior? Here, we address these questions by investigating the representations formed by LLMs during an in-context concept inference task. We found that LLMs can flexibly derive concepts from linguistic descriptions in relation to contextual cues about other concepts. The derived representations converge toward a shared, context-independent structure, and alignment with this structure reliably predicts model performance across various understanding and reasoning tasks. Moreover, the convergent representations effectively capture human behavioral judgments and closely align with neural activity patterns in the human brain, providing evidence for biological plausibility. Together, these findings establish that structured, human-like conceptual representations can emerge purely from language prediction without real-world grounding, highlighting the role of conceptual structure in understanding intelligent behavior. More broadly, our work suggests that LLMs offer a tangible window into the nature of human concepts and lays the groundwork for advancing alignment between artificial and human intelligence.
[…]
Significance Statement
Large language models (LLMs) show intriguing human-like behaviors despite being trained solely via language prediction. Are these models developing human-like concepts central to human understanding? Here, we demonstrate that LLMs can flexibly derive concepts from linguistic descriptions across varying contexts. The derived representations converge toward a shared, context-independent structure, which effectively predicts human behavioral judgments and aligns closely with neural activity patterns in the human brain. Our findings reveal that critical aspects of human concepts are learnable purely from language prediction. Rather than relying on real-world grounding, LLMs organize concepts through meaningful interrelationships preserved across contexts, providing a tractable window into human conceptual representation and organization.
@@ 1 Introduction
Humans are able to construct mental models of the world and use them to understand and navigate their environment [1, 2]. Central to this ability is the capacity to form broad concepts that constitute the building blocks of these models [3]. These concepts, often regarded as mental representations, capture regularities in the world while abstracting away extraneous details, enabling flexible generalization to novel situations [4]. For example, the concept of sun can be instantly formed and deployed across diverse contexts: observing it rise or set in the sky, yearning for its warmth on a chilly winter day, or encountering someone who exudes positivity. Its role in the solar system can be analogized to the nucleus in an atom, enriching learning and understanding. The nature of concepts has long been a focus of inquiry across philosophy, cognitive science, neuroscience, and linguistics [5, 6, 7, 8, 9, 10, 11]. These investigations have uncovered diverse properties that concepts need to satisfy, often framed by the long-standing divide between symbolism and connectionism. Symbolism emphasizes discrete, explicit symbols with structured and compositional properties, enabling abstract reasoning and recombination of ideas [12, 13]. In contrast, connectionism conceptualizes concepts as distributed, emergent patterns across networks, prioritizing continuity and gradedness, which excel in handling noisy inputs and learning from experience [14, 15]. Although there is growing consensus on the need to integrate the strengths of both paradigms to account for the complexity and richness of human concepts [16, 17], reconciling the competing demands remains a significant challenge.
@@
Recent advances in large language models (LLMs) within artificial intelligence (AI) have exhibited human-like behavior across various cognitive and linguistic tasks, from language generation [18, 19] to decision-making [20] and reasoning [21, 22, 23]. It has sparked intense interest and debate about whether these models are approaching human-like cognitive capacities based on concepts [24, 25, 26, 27, 28, 29]. Some argue that LLMs, trained solely on text data for next-token prediction, operate only on statistical associations between linguistic forms and lack concept-based understanding grounded in physical and social situations [30, 31, 28]. This argument is evidenced by their inconsistent performance, which often reveals vulnerabilities such as non-human-like errors and over-sensitivity to minor contextual variations [20, 32, 33]. Conversely, others contend that the performance of a system alone is insufficient to characterize its underlying competence [34], and the extent to which concepts should be grounded remains open [35, 36]. Instead, language may provide essential cues for people’s use of concepts and enable LLMs to approximate fundamental aspects of human cognition, where meaning arises from the relationships among concepts [37, 38]. Despite the conflicting views, there is broad consensus on the central role of concepts in understanding both human and machine cognition. The core questions driving the debate are whether LLMs possess concepts that embody the key properties of human concepts, and how such concepts are represented, organized, and linked to behavior.
@@
To address these questions, we investigated conceptual representations emerging from language prediction in LLMs. Our approach unfolded in three stages. First, we examined LLMs’ capacities to form and organize conceptual representations, focusing on the definitional and structured properties of human concepts. We guided LLMs to derive concepts from descriptions with a few contextual demonstrations (Fig. 1). The models’ outputs offered a behavioral lens to evaluate the derived representations. We then explored the organization of these representations by analyzing how their relational structure varied across contexts, and further examined how this structure related to model performance on a range of language understanding and reasoning tasks. Second, we assessed the extent to which the LLM-derived representational structure aligns with psychological measures and captures rich, context-dependent human knowledge. These properties were tested by using computations over the representations to predict human behavioral judgments. Finally, we examined the biological plausibility of these representations by mapping them to neural activity patterns in the human brain. Our experiments spanned thousands of naturalistic object concepts [39] and beyond, providing a comprehensive analysis of LLMs’ conceptual representations. The findings reveal that language prediction alone can give rise to human-like conceptual representations in LLMs. These representations are structured in a way that integrates critical aspects of human concepts, combining the definitional and structural focus of symbolism with the continuity and graded nature of connectionist models. Alignment with this structure reliably predicts model performance across diverse tasks. Collectively, this work underscores the profound connection between language and human conceptual structure, and provides a simple, generalizable framework for deriving conceptual representations from LLMs as a reusable resource, opening new pathways for exploring the nature of human concepts.
Figure 1: Illustration of the reverse dictionary task as a conceptual probe. A Transformer-based LLM is presented with NN description–word pairs as demonstrations in context, followed by a query description. The model is then prompted to encode the query description into a conceptual representation and predict the term that best matches the described concept.
@@ 2.1 Reverse dictionary as a conceptual probe
We reframed the reverse dictionary task as a domain-general probe of LLMs’ capacity for concept inference. People can identify and think about concepts from definitions or descriptions, even when these are incomplete or imprecise [3]. The reverse dictionary task provides a natural simulation of this process. Formally, given a description of a concept $c$, a system encodes it into a latent representation conditioned on context, $\mathbf{h}_{c}=\textrm{encode}\left(\textrm{description}\mid\textrm{context}\right)$, and generates a corresponding term: $\textrm{predict}\left(\mathbf{h}_{c}\right)$. Rather than relying on a single, potentially ambiguous word (e.g., “bat”), this process requires abstraction beyond surface forms, unifying diverse linguistic expressions of the same underlying concept (e.g., “a round, glowing object that hangs in the night sky” and “Earth’s natural satellite” both correspond to the concept moon). The generated term (e.g., “moon”) provides a tangible lens on whether the inferred representation aligns with the intended concept. This paradigm thus offers a targeted probe into a system’s capacity to contextually form consistent and shareable representations, capturing two key properties of human concepts: definitions and language generation.
To guide LLMs through the process, we took advantage of their in-context learning ability and presented them with a small number of demonstrations ($N$) in a reverse-dictionary format ($\textrm{Description}]\Rightarrow[\textrm{Word}]$”), followed by a query description (Fig. 1). We then compared model-generated completion to the intended term of the query concept, thereby evaluating how well the model-derived conceptual representations aligned with human understanding. Importantly, the demonstrations provided minimal and controllable context, and the query description served as an input stimulus for concept inference. We can analyze how models construct conceptual representations in relation to contextual cues from other concepts by varying the number of demonstrations and the description–word pairings within them.
@@ 2.2 Deriving concepts from definitional descriptions
We investigated whether LLMs can construct concepts from definitional descriptions through the reverse dictionary task. Leveraging data from the THINGS database [40], we prompted LLMs with randomly selected description–word pair demonstrations and examined their capacity to predict appropriate terms for the query descriptions (Materials and Methods). Fig. 2 shows the performance of LLaMA3-70B, a state-of-the-art open-source LLM, averaged across five independent runs at varying numbers of demonstrations. The exact match accuracy improved progressively as the number of demonstrations increased from 1 to 24, rising from $79.51\%$ ($\pm 0.30\%$) to $89.45\%$ ($\pm 0.30\%$). Gains were marginal beyond this threshold. These results indicate that, with a few demonstrations as context, LLMs can effectively combine words into coherent representations and reliably infer the corresponding concepts, despite varying contextual cues.
Figure 2: Performance of LLaMA3-70B on the reverse dictionary task measured through exact match accuracy. Black lines show performance when the model is given NN correct demonstrations sampled from the training set and evaluated on an independent test set. Blue and red lines show performance when one misleading demonstration—a description paired with a proxy label (capital letter, random string, or random word, as specified in the figure titles)—is added to N1N-1 correct demonstrations of other concepts. The model is queried with the description identical to the misleading one. The blue line shows the frequency with which the model reproduces the proxy label, while the red line indicates how often it generates the correct word for the query concept given contextual information from other concepts. Shaded areas denote $95\%$ bootstrapped confidence intervals, calculated from 10,000 resamples over five independent runs.
@@
Furthermore, we extended our analyses to a broader range of descriptions and concepts to assess the generality of our findings derived from the THINGS database. Using data from WordNet, our extended analysis demonstrated the strong adaptability of LLMs across concepts differing in concreteness and word classes (Fig. S8A). Consistent results were also observed across various descriptions of the same concept (Fig. S8B). Introducing varying degrees of word order permutations to the query descriptions revealed that LLM predictions were sensitive to linguistic structure degradation when combining words to form concepts (Fig. S8C). These findings highlight the model’s effectiveness in concept inference and its ability to capture at least some of the compositional structure in natural language. Beyond LLaMA3-70B, we tested 66 additional open-source LLMs with different model architectures, scales and training data. The results revealed trends similar to those observed with LLaMA3-70B. They also indicated that larger models generally perform better and more effectively leverage contextual cues for concept inference (SI Appendix, section SI 1.A, Fig. S8 DF).
@@ 2.3 Uncovering a context-independent conceptual structure
Figure 3: LLMs converge toward a similar representational structure of concepts. (A) Pairwise alignment (RSA) of LLaMA3-70B conceptual representations across different contextual demonstrations. Axes indicate the number of demonstrations, and each cell shows the alignment from a single run. (B) LLM performance on the reverse dictionary task reflects alignment with the representations formed with 120 demonstrations. Each point corresponds to a representation space formed with NN demonstrations, with the x-axis indicating exact match accuracy and the y-axis denoting alignment with the 120-demonstration space. Error bars indicate $95\%$ confidence intervals from 10,000 bootstrap resamples across five runs. (C) A t-SNE visualization of conceptual representations formed by different LLMs with 24 demonstrations. Distances are calculated as $1-\textrm{alignment}$, averaged over five runs. Each point corresponds to an LLM, plotted in proportion to its complexity and color-coded by exact match accuracy on the reverse dictionary task. Better-performing models (blue) exhibit more aligned representations. (D) Alignment with LLaMA3-70B conceptual representations predicts model performance across downstream tasks. The x-axis shows the alignment with LLaMA3-70B representations, and the y-axis indicates overall performance, computed as accuracy averaged across a range of language understanding and reasoning tasks. Each point represents a different model and is color-coded by model complexity. Linear fits are shown as straight lines, with shaded areas representing $95\%$ confidence intervals derived from 10,000 bootstrap resamples. Overall, models with higher complexity align more closely with the LLaMA3-70B. An exception is Qwen2-0.5B, which shows relatively high complexity but low alignment, possibly due to constraints in model scale or training data quality.
To uncover how concepts are represented and organized within LLMs, we looked into the representation spaces they constructed for concept inference and analyzed the interrelationships among the conceptual representations. We characterized the representation spaces formed under different contexts by their relational structure, captured through a (dis)similarity matrix, and measured pairwise alignment by correlating these matrices [41, 42] (Materials and Methods). For concepts in the THINGS database, the alignment between representation spaces gradually increased as the number of demonstrations rose from 1 to 24, with diminishing gains beyond this threshold (Fig. 3A). The alignment with the space formed by 120 demonstrations increased from $0.800$ ($\pm 0.017$) with one demonstration to $0.970$ ($\pm 0.003$) after 24 demonstrations. This alignment demonstrated a strong correlation with the model’s exact match accuracy on concept inference ($\rho=0.976$, $P<0.0001$, $95\%$ CI: 0.7300.7301.0001.000) (Fig. 3B). These results suggest that LLMs are able to construct a coherent and context-independent relational structure, which is reflected in their concept inference capacity. Additional metrics corroborated this conclusion (SI Appendix, section SI 1.B), highlighting that intricate relationships among concepts are consistently preserved within LLMs’ representation spaces across varying contexts. This invariance supports the generalization of knowledge encoded in these relationships.
To examine whether different LLMs trained for language prediction develop similar conceptual representations, we compared the representation spaces they formed based on 24 demonstrations. Using t-SNE [43], we visualized pairwise alignments between models and observed that LLMs with over $70\%$ exact match accuracy on concept inference clustered closely, whereas those with accuracy below $50\%$ exhibited greater dispersion (Fig. 3C). This indicates that better-performing models share more similar relational structure of concepts, while those with weak performance diverge in distinct ways.
We further conducted a correlational analysis to assess how the relational structure relates to model behavior more broadly. Models whose structures aligned more closely with that of LLaMA3-70B achieved higher accuracy not only on the reverse dictionary task ($\rho=0.870$, $P<0.0001$, $95\%$ CI: 0.7560.7560.9400.940) but also across a diverse set of language understanding and reasoning tasks ($\rho=0.831$, $P<0.0001$, $95\%$ CI: 0.6990.6990.9160.916; Fig. 3D; see SI Appendix, Fig. S10 for results on each task). Additionally, when quantifying model complexity based on scale and training data (Materials and Methods), we found that higher-complexity LLMs tended to align better with LLaMA3-70B, though exceptions likely stemmed from constraints in model scale or training data quality (Fig. 3D). These findings suggest that, with sufficient model scale and extensive, high-quality training data, LLMs can converge toward a shared conceptual structure. This convergence is predictive of their performance on understanding and reasoning tasks, and arises naturally from the objective of language prediction, without requiring real-world reference.